SPECTROSCOPIC STUDY OF CYANOACETYLENE CATION: SLOW PHOTO-ELECTRON SPEC-TROSCOPY AND AB-INITIO INVESTIGATIONS

\underline{AHMED MAHJOUB}, MARTIN SCHWELL, YVES BENILAN, NICOLAS FRAY and MARIE-CLAIRE GAZEAU, LISA UMR CNRS 7583, Univ. Paris Est Creteil & Univ. Paris Diderot, Institut Pierre Simon Laplace,61 Ave du General de Gaulle, 94010, Creteil, France; GUSTAVO A. GARCIA and FRANCOIS GAIE-LEVERL, Synchrotron SOLEIL, LOrme des Merisiers, St.Aubin, B.P. 48, 91192, Gif-sur-Yvette Cedex, France; NORBERT CHAMPION and SYDNEY LEACH, LERMA UMR CNRS 8112, Observatoire de Paris-Meudon, 5 place Jules-Jansen, 92195, Meudon, France.

Cyanoacetylene is one of the key minor constituents in the atmosphere of Titan. The ion HCCCN+ has been detected in this atmosphere and it is supposed to be formed by the reaction between C3H2+ and atomic nitrogen^{a}. We present here a spectroscopic investigation of the cyanoacetylene cation using photoexcitation of the neutral by vacuum-ultraviolet (VUV) synchrotron radiation coupled to a velocity map imaging electron/ion coincidence spectrometer^{b}. The cation spectroscopy is studied by the Slow Photoelectron Spectroscopy technique (SPES) (figure below) and the Total Ion Yield (TIY). The TIY has been calibrated to absolute units using the known propane absolute cross-section. Quantum chemical calculations are performed in order to interpret these spectra. These calculations deal with the equilibrium geometries, electronic-state patterns and evolutions, and harmonic and anharmonic wavenumbers. Through this study, we observe, in the auto-ionization region above the ionization energy, a number of Rydberg series of neutral cyanoacetylene. These Rydberg series converge to the first and second excited states of the cation. Acknowledgments. We are indebted to the general technical staff of Synchrotron Soleil for the running facility.

aC. Barrientos, P. Redondo and A. Largo J. Chem. Phys. A \underline{104}(49), 11541-11548. 2000

bL. Nahon, N. De Oliveria,J. F. Gil,B. Pilette,O. Marcouill, B. La garde and F. Polack Journal of Synchrotron Radiation \underline{19}(4), 508-520; 2012